Use dimensional analysis to express the wave velocity in terms of the tension in the string and mass per unit length of the string.
1
Expert's answer
2019-05-20T13:47:44-0400

Let the wave velocity is dependent on the tension and mass per unit length. Then


"v=F^{\alpha}\mu^{\beta}"

Since dimensions of velocity, force and mass per unit length are as follows,


"[v]=LT^{-1}"

"[F]=MLT^{-2}"

"[\mu]=ML^{-1}"

we obtain

"LT^{-1}=\left(MLT^{-2}\right)^{\alpha}(ML^{-1})^{\beta}"

"LT^{-1}=M^{\alpha+\beta}L^{\alpha-\beta}T^{-2\alpha}"

So


"\left\{\begin{matrix}
 \alpha+\beta &=& 0 \\
 \alpha-\beta &=& 1\\
-2\alpha &=& -1
\end{matrix}\right."

Solution

"\left\{\begin{matrix}
 \alpha &=& 1/2\\
 \beta &=& -1/2
\end{matrix}\right."

Finally

"v=F^{1/2}\mu^{-1/2}=\sqrt{\frac{F}{\mu}}"